The most recent financial statements for Minnie's Manufacturing Co. are shown below:
Income Statement Balance Sheet
  Sales $ 96,000   Current assets $ 36,000   Debt $ 46,800
  Costs 68,550   Fixed assets 95,600   Equity 84,800
  Taxable income $ 27,450      Total $ 131,600      Total $ 131,600  
  Tax 6,588
  Net Income $ 20,862

Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 35 percent dividend payout ratio. No external financing is possible.

What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

ROA=Net income/Total assets

=20,862/131,600

=0.158525836

Retention ratio=1-payout ratio

=(1-0.35)=0.65

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.158525836*0.65)/[1-(0.158525836*0.65)

=11.49%(Approx).